OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 16x = 3

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

are you talking bout makin it a diff of squares

OpenStudy (anonymous):

what number completes the square ?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

because x^2+2*8*x+8^2-8^2=3 (x+8)^2-64=3

OpenStudy (anonymous):

so is it 3 or 64?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

What is (x+4)^2?

OpenStudy (anonymous):

why you asking on my question?

OpenStudy (anonymous):

x^2+2*4*x+16

OpenStudy (anonymous):

It is the answer to your question....

OpenStudy (anonymous):

it cant be.

OpenStudy (anonymous):

boy?

OpenStudy (anonymous):

Ok, (x+8)^2....

OpenStudy (anonymous):

@estudier is right if can solve the simple one then you can solve your own problems.

OpenStudy (anonymous):

Completing the square step by step x^2 + 16x = 3 Step 1 Choose b=16 where b= coefficient of x in this case b=16 Step 2 add ( b/2)^2 to both sides , b/2 = 16/2 =8 so (b/2)^ = 64 this response your answer X^2 + 16x + 64 = 64+ 3=67 (x+ 8)^2 = 67 So you must add (b/2)^2 = 64

OpenStudy (anonymous):

\[x=\sqrt(67)-8 \]

OpenStudy (anonymous):

The person were asking for a number need to for completing the square not for the solution of x ...